# ZKP-FedEval: Verifiable and Privacy-Preserving Federated Evaluation using Zero-Knowledge Proofs

### v1

@@ I Introduction
Federated Learning (FL) has emerged as a practical method for training machine learning models on decentralized data, offering privacy advantages over centralized approaches [1]. Nonetheless, the evaluation phase in FL poses privacy challenges. Even though clients compute scalar metrics such as loss or accuracy on local data, sharing these values may enable inference attacks that leak sensitive information [2, 3]. Existing countermeasures such as Differential Privacy [4] introduce noise that degrades model utility, and secure aggregation methods [5] do not verify the computation’s integrity.
@@ II Related Work
Recent efforts have explored the application of ZKPs in machine learning [10, 11] and federated learning [12, 13]. In contrast, our approach focuses on the evaluation phase using a lightweight, threshold-based ZKP protocol implemented entirely in-house without external dependencies.

### v2

@@ I Introduction
Federated Learning (FL) has emerged as a practical method for training machine learning models on decentralized data, offering privacy advantages over centralized approaches [1]. However, its distributed nature introduces significant trust and security challenges throughout its lifecycle. For instance, the training phase is vulnerable to data poisoning attacks, which requires robust defense mechanisms to ensure client integrity [2]. Orthogonal to securing training, the evaluation phase poses its own privacy challenges. Sharing even scalar metrics like loss or accuracy may enable inference attacks that leak sensitive information [3, 4]. Existing countermeasures such as Differential Privacy [5] introduce noise that degrades model utility, and secure aggregation methods [6] do not verify the computation’s integrity.
@@ II Related Work
Recent efforts have explored the application of ZKPs in machine learning [12, 13] and federated learning [14, 15]. The versatility of ZKPs allows them to secure various system layers, from proving hardware identity in blockchain-based IoT systems [16] to verifying complex computations. In contrast to prior work in FL, our approach focuses specifically on the evaluation phase, using a lightweight, threshold-based ZKP protocol implemented entirely in-house without external dependencies.